### v1

@@ Abstract
In this paper, we investigate the potential of image-to-image translation (I2I) techniques for transferring realism to 3D-rendered facial images in the context of Face Recognition (FR) systems. The primary motivation for using 3D-rendered facial images lies in their ability to circumvent the challenges associated with collecting large real face datasets for training FR systems. These images are generated entirely by 3D rendering engines, facilitating the generation of synthetic identities. However, it has been observed that FR systems trained on such synthetic datasets underperform when compared to those trained on real datasets, on various FR benchmarks. In this work, we demonstrate that by transferring the realism to 3D-rendered images (i.e., making the 3D-rendered images look more real), we can boost the performance of FR systems trained on these more photorealistic images. This improvement is evident when these systems are evaluated against FR benchmarks utilizing real-world data, thereby paving new pathways for employing synthetic data in real-world applications  [^1] The source code will be made available and reproducible upon acceptance at here..
@@ 1 Introduction
It is difficult to quantify the benefits gained through synthetic datasets, as they often fall short of the performance achieved by the pre-trained FR systems used during their generation phase.
@@ 4 Experiments
For training unpaired I2I methods, specifically, VSAIT, UNSB, and DECENT, we randomly selected five shards for the source domain (i.e., 3D-rendered human face images), each containing $20,000$ images from the DigiFace1M dataset. Similarly, for the target domain, we randomly selected five shards, each containing $20,000$ images from the FFHQ dataset, and experimented with training these three models using multiple combinations of source and target shards. After training the realism transfer methods, we selected two according to the time they needed to process an image and qualitative examination of the processed images, which are depicted by Time/Image (s) and Qualitative Ex respectively in Table 1. The processing time was measured on an NVIDIA RTX 3090 Ti across all methods. Figure 3 presents some qualitative results of various methods. As can be qualitatively observed from Figure 3, CodeFormer generally performed very well across all samples, preserving the entire facial structure. In contrast, VSAIT, DECENT, and UNSB did not consistently produce quality images. Notably, these models sometimes dislocated parts of the images, resulting in multiple eyes and mouths. Surprisingly, as we will demonstrate in the next section, VSAIT significantly boosted the performance of FR systems. Here, ‘UNSB-NE-1’ and ‘UNSB-NE-5’ refer to the number of Neural Evaluation (NE) steps of the method; for more details, please refer to the original paper. Finally, images produced by DDIM-Inversion appear as smoothed-out versions of the originals. Among the examined methods, we chose CodeFormer because of its good quality and reasonable compute time and VSAIT for its lower compute time and slightly better quality than other I2I methods for the final FR experiments in the next section.
@@ 4.1 Face Recognition Experiments
For a fair comparison between different methods, we trained an FR system consisting of a ResNet50 backbone as modified in ArcFace’s implementation [11], with the AdaFace [21] head for contrastive loss. We trained a separate network for each of the methods mentioned in the previous section, namely, the original DigiFace1M, and translated versions of the images generated using CodeFormer and VSAIT. We also included an FR baseline that methods like DCFace and IDiffFace are using; we used the pre-trained model provided by the AdaFace paper, which was trained on the WebFace4M dataset. For FR benchmarking, we considered various datasets including LFW [16], CFPFF [38], CPLFW [52], CALFW [53], AgeDB [28], which consist of high-quality images with various lighting, poses, and ages. We also benchmarked against IJB-C [26], which is amongst the most challenging FR benchmarks in the literature. The results are reported in Table 2 and Table 3.
@@
Table 2: Results of different synthetic data generation methodologies used to train multiple FR systems evaluated on the LFW, CPFF, CPLFW, CALFW, and AgeDB, the last column is the average test accuracy over these five datasets. We are reporting mean and std over multiple runs of experiments in each row.
Transfer Method Dataset Type SynGen Req LFW CFPFF CPLFW CALFW AGEDB Avg
@@
In the case of Table 2, compared to the model trained on the DigiFace1M, we observed an average improvement of $2.0\%$ over the FR model trained on images generated after transferring them using CodeFormer and VSAIT, with CodeFormer demonstrating a slight advantage in all datasets: LFW, CFPFF, CPLFW, CALFW, and AgeDB. However, it can be observed that there is a significant performance gap between the model trained on the WebFace4M and all other methods, including the DCFace and IDiffFace models, which use such a competitive FR system for generating their images.
@@ 5 Conclusion and Future Work
Surprisingly, we found that by solely employing transfer models that do not incorporate identity labels in their training paradigm, we can boost the performance of FR systems across all benchmarks—LFW, CFPFF, CPLFW, CALFW, AGEDB, and IJB-C—by $2\%$ to $5\%$. This improvement is observed in comparison with models trained on the original DigiFace1M dataset, thereby narrowing the performance gap with models that use pre-trained FR data for generating their data. Moreover, this approach moves us closer to our ultimate goal of achieving performance parity with models trained on real data. This opens new avenues for exploring the use of transfer methodologies in the domain of data enhancement for improved downstream model performance.
As mentioned previously, this work aimed to demonstrate the potential for realism transfer in the domain of Face Recognition systems, paving the way for the development of more sophisticated transfer methods for this task. Given that the pipeline for developing a new transfer method and applying it to the entirety of a source dataset is cumbersome and time-consuming—especially since it necessitates multiple trainings of the FR system on the generated data for conclusions—a future work, could be to explore a form of quality assessment metric. This metric would correlate with the final performance of the FR system when trained on the generated dataset, allowing for the evaluation of the transferred data independently. This approach could significantly streamline the process of assessing the potential of newly generated datasets for FR applications.
@@ Acknowledgment
This research is based on work conducted in the SAFER project and supported by the Hasler Foundation’s Responsible AI program. We also thank Dr. Damien Teney for his thoughtful feedback.

### v2

@@ Abstract
[…]
@@ 1 Introduction
It is difficult to quantify the benefits gained through synthetic datasets, as they often fall short of the performance achievesd by the pre-trained FR systems used during their generation phase.
@@ 4 Experiments
For training unpaired I2I methods, specifically, VSAIT, UNSB, and DECENT, we randomly selected five shards for the source domain (i.e., 3D-rendered human face images), each containing $20,000$ images from the DigiFace1M dataset. Similarly, for the target domain, we randomly selected five shards, each containing $20,000$ images from the FFHQ dataset, and experimented with training these three models using multiple combinations of source and target shards. After training the realism transfer methods, we selected two according to the time they needed to process an image and qualitative examination of the processed images, which are depicted by Time/Image (s) and Qualitative Ex respectively in Table 1. The processing time was measured on an NVIDIA RTX 3090 Ti across all methods. Figure 3 presents some qualitative results of various methods. As can be qualitatively observed from Figure 3, CodeFormer generally performed very well across all samples, preserving the entire facial structure. In contrast, VSAIT, DECENT, and UNSB did not consistently produce quality images. Notably, these models sometimes dislocated parts of the images, resulting in multiple eyes and mouths. Surprisingly, as we will demonstrate in the next section, VSAIT boosted the performance of FR systems. Here, ‘UNSB-NE-1’ and ‘UNSB-NE-5’ refer to the number of Neural Evaluation (NE) steps of the method; for more details, please refer to the original paper. Finally, images produced by DDIM-Inversion appear as smoothed-out versions of the originals. Among the examined methods, we chose CodeFormer because of its good quality and reasonable compute time and VSAIT for its lower compute time and slightly better quality than other I2I methods for the final FR experiments in the next section.
@@ 4.1 Face Recognition Experiments
For a fair comparison between different methods, we trained an FR system consisting of a ResNet50 backbone as modified in ArcFace’s implementation [11], with the AdaFace [21] head for contrastive loss. We trained a separate network for each of the methods mentioned in the previous section, namely, the original DigiFace1M, and translated versions of the images generated using CodeFormer and VSAIT. We name the translated dataset RealDigiFace. We also included an FR baseline that methods like DCFace and IDiffFace are using; we used the pre-trained model provided by the AdaFace paper, which was trained on the WebFace4M dataset. For FR benchmarking, we considered various datasets including LFW [16], CFPFP [38], CPLFW [52], CALFW [53], AgeDB [28], which consist of high-quality images with various lighting, poses, and ages. We also benchmarked against IJB-C [26], which is amongst the most challenging FR benchmarks in the literature. The results are reported in Table 2 and Table 3.
@@
Table 2: Results of different synthetic data generation methodologies used to train multiple FR systems evaluated on the LFW, CFPFP, CPLFW, CALFW, and AgeDB, the last column is the average test accuracy over these five datasets. We are reporting mean and std over multiple runs of experiments in each row.
Transfer Method Dataset Type SynGen Req LFW CFPFP CPLFW CALFW AGEDB Avg
@@
In the case of Table 2, compared to the model trained on the DigiFace1M, we observed an average improvement of $2.0\%$ over the FR model trained on images generated after transferring them using CodeFormer and VSAIT, with CodeFormer demonstrating a slight advantage in all datasets: LFW, CFPFP, CPLFW, CALFW, and AgeDB. However, it can be observed that there is a significant performance gap between the model trained on the WebFace4M and all other methods, including the DCFace and IDiffFace models, which use such a competitive FR system for generating their images.
@@ 5 Conclusion and Future Work
Surprisingly, we found that by solely employing transfer models that do not incorporate identity labels in their training paradigm, we can boost the performance of FR systems across all benchmarks—LFW, CFPFF, CPLFW, CALFW, AGEDB, and IJB-C—by $2\%$ to $5\%$. This improvement is observed in comparison with models trained on the original DigiFace1M dataset, thereby narrowing the performance gap with models that use pre-trained FR data for generating their data. Moreover, this approach moves us closer to our ultimate goal of achieving performance parity with models trained on real data. This opens new avenues for exploring the use of transfer methodologies in the domain of data enhancement for improved downstream model performance. Given that the pipeline for developing a new transfer method and applying it to the entirety of a source dataset is cumbersome and time-consuming—especially since it necessitates multiple trainings of the FR system on the generated data for conclusions—a future work, could be to explore a form of quality assessment metric. This metric would correlate with the final performance of the FR system when trained on the generated dataset, allowing for the evaluation of the transferred data independently. This approach could significantly streamline the process of assessing the potential of newly generated datasets for FR applications.
@@ Acknowledgment
This research is based on work conducted in the SAFER project and supported by the Hasler Foundation’s Responsible AI program. We would also like to extend our gratitude to Dr. Damien Teney for his thoughtful feedback.